Patriot Co. manufactures and sells three products: red, white, and blue. Their unit selling prices are red, $61; white, $91; and blue, $116. The per unit variable costs to manufacture and sell these products are red, $46; white, $66; and blue, $86. Their sales mix is reflected in a ratio of 5:4:2 (red:white:blue). Annual fixed costs shared by all three products are $156,000. One type of raw material has been used to manufacture all three products. The company has developed a new material of equal quality for less cost. The new material would reduce variable costs per unit as follows: red, by $9; white, by $19; and blue, by $9. However, the new material requires new equipment, which will increase annual fixed costs by $26,000.

1. Assume if the company continues to use the old material, determine its break-even point in both sales units and sales dollars of each individual product. (Round composite units up to next whole number.)
Answer #1

X fx 1 | B I ratio red white selling price per unit $61 $91 $ 116 total per composite unit $ 305 $364 $ 232 $ 901 blue total

for formulas and calculations, refer to the image below -

B2 | x 1 fc D B A 1 ratio I selling price per unit 1 red white blue 5 4 2 total per composite unit =D3*E3 =D4*E4 =D5*E5 =SUM(

In case you have any query, kindly ask in comments.